OpenStudy (anonymous):

express 3/5 meters in centimeters

OpenStudy (anonymous):

A centimeter is 1/100 th of a meter. To find out how many are in x meters divide x by 1/100. In this case. \[\frac{3/5}{\frac{1}{100}}=60\]

OpenStudy (anonymous):

3/5=3/5*100=3*20=60